OpenStudy (anonymous):

Help me check my work? Write the equation of the line that goes through the points (-1, 3) and (4, -7) in general form. Also, could you clearly show and explain each step, thanks in advance:)

OpenStudy (anonymous):

general form: y=mx+b equation for slope of a line: m=rise/run m=(y2-y1)/(x2-x1) m=(-7-3)/(4-(-1))=-2 so.. y=-2x+b now plug in one of the given points to solve for b, I chose (4,-7) -7=-2(4)+b b=1 so equation of the line is: y=-2x+1

OpenStudy (anonymous):

I tried it, and just wanted to make sure it was correct? Since they asked for the answer to be in general form (which is standard form) of a line: Ax + By = C where A is a positive integer, and B, and C are integers First I found the slope between the two points: m = (y2-y1)/(x2-x1) m = (4-3)/(4-1) m = 1/3 (slope) Then I used one of the two points (1,3) and the slope (1/3) and plug it into the point-slope form: y-y1 = m(x-x1) y-3 = (1/3)(x-1) multiplying both sides by 3: 3y-9 = x-1 subtracting x from both sides: -x + 3y-9 = -1 adding 9 to both sides: -x + 3y = 8 is that correct? it seems pretty correct to me

OpenStudy (anonymous):

it looks correct

OpenStudy (anonymous):

general form of a line is y=mx+b if you want it in Ax+By=C form then it would be y+2x=1 from the work I did above

OpenStudy (anonymous):

oh wait. general form is Ax + By + C = 0 so would it then be -x + 3y = 8 (+0) -x + 3y - 8 = 0 ?

OpenStudy (anonymous):

I think you are using the wrong points for you y2, y1, x2, x1 y2=-7 y1=3 x2=4 x1=-1 use those and try to solve it again. With the form you want the equation 2x+y-1=0